OpenStudy (anonymous):

Find the solution of this system of equations. separate the x-and y- values with a comma. x=8+y -2x+6y=-20 (I think I messed up when trying to solve.)

OpenStudy (anonymous):

x=8+y -2x+6y=-20 ?

OpenStudy (anonymous):

Problem: Solve x=8+y;−2x+6y=−20 Steps: I will try to solve your system of equations. x=y+8;−2x+6y=−20 Step: Solve x=y+8 for x: Step: Substitute (y+8) for x in −2x+6y=−20: −2x+6y=−20 −2(y+8)+6y=−20 4y−16=−20(Simplify both sides of the equation) 4y−16+16=−20+16(Add 16 to both sides) 4y=−4 4y4=−44(Divide both sides by 4) y=−1 Step: Substitute (−1) for y in x=y+8: x=y+8 x=−1+8 x=7(Simplify both sides of the equation) Answer: x=7 and y=−1

OpenStudy (anonymous):

yes and thank you

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

wait it says that it is wrong.

OpenStudy (anonymous):

it shouldn't be hold on let me re calculate

OpenStudy (anonymous):

are you graphing it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

my calculations where right i just went over them

OpenStudy (anonymous):

I am not sure then but its okay thanks anyways.